A girl went to school & the teacher asked how old are you? She replied I don't know, bt I know that I'm half my mother's age. The teacher asked ok how old is your mother? She said I don't know but I know that my Dad is 5 yrs older than my mother. Teacher, how old is yor dad? She said I don't know but I know that if we add our ages the answer is 100 yrs... Nw how old is the girl, her mum & her dad?

Answer by lwsshak3(11628)

A girl went to school & the teacher asked how old are you? She replied I don't know, bt I know that I'm half my mother's age. The teacher asked ok how old is your mother? She said I don't know but I know that my Dad is 5 yrs older than my mother. Teacher, how old is yor dad? She said I don't know but I know that if we add our ages the answer is 100 yrs... Nw how old is the girl, her mum & her dad?
***
let x=mum's age
x/2=girl's age
x+5=dad's age
..
x+x/2+x+5=100
lcd:2
2x+x+2x+10=200
5x=190
x=38
x/2=19
x+5=43
mum's age=38
girl's age=19
dad's age=43